# The Fate of Entanglement

### v2

@@ Abstract
Icosahedral molecule.— We illustrate the above results with a simple quantum system: the anti-ferromagnetic Ising model on the 12-spin icosahedron, see Fig. 2a, with Hamiltonian $H=J\sum_{{\rm bonds}\,\langle i,j\rangle}\sigma_{i}^{x}\sigma_{j}^{x}-h\sum_{{\rm sites}\,i}\sigma_{i}^{z}$. The first term corresponds to an anti-ferromagnetic interaction $J>0$ that tends to anti-align neighbouring spins along $x$, while the second one is a transverse field that polarises all spins along $z$. Such a molecular quantum magnet is a combination of identical triangular faces, and thus possesses strong geometric frustration. In what follows we shall measure energy in units of the exchange coupling by setting $J=1$. We investigate the fate of 2- and 3-spin entanglement as a function of magnetic field, temperature, time a separation by considering three distinct quantities. See the Supplemental Material (SM) for their precise definitions. First, we compute a measure of 2-party entanglement, focusing on a pair of adjacent spins, called the logarithmic negativity $\mathcal{E}$ [30, 31]. In the case of two spins, $\mathcal{E}=0$ implies that the reduced density matrix is separable [32], whereas for entangled states we have $\mathcal{E}>0$. Second, we study the 3-spin entanglement on a triangular face of the molecule via the geometric entanglement $\mathcal{D}$ [3] defined as the smallest distance between the state and the separable continent,
$\mathcal{D}=\min_{\rho_{\rm sep}}d(\rho,\rho_{\rm sep}),$ (2)
where the minimisation is over all $\rho_{\rm sep}$ living on the continent. This powerful measure detects all forms of entanglement. Finally, we employ a criterion $W$ that indicates the presence of genuine 3-party entanglement when $W>0$ [33]. This type of entanglement relates all three parties, and cannot be described in term 2-party entanglement. When $W=0$, a definitive conclusion cannot be made. We note that both $\mathcal{D}$ and $W$ require a minimisation procedure in a multi-parameter space, making their evaluation more demanding than simpler measures such as $\mathcal{E}$. As a prerequisite, we checked that the FRH holds for subregions of adjacent spins with 1, 2 and 3 sites by obtaining the $2^{12}$ eigenstates via exact numerical diagonalization, see SM.
@@
We first consider the fate of entanglement with varying magnetic field at zero temperature, so that the evolution is parameterized by $s=h$, see Fig. 2b. At $h\!=\!0$, the 2-spin reduced density matrix is a full-rank separable state, implying the existence of a value $h_{2}^{*}>0$ of the field below which the state is separable, and hence $\mathcal{E}(h<h_{2}^{*})=0$. We find $h_{2}^{*}\approx 0.6$. In contrast, the 3-spin reduced density matrix is not separable at small $h$ and we have $\mathcal{D}>0$. In the opposite limit of $h\to\infty$, the system becomes fully polarized and is in a pure product state. Both 2- and 3-spins density matrices are thus pure product states, i.e., of rank one, and lie on the shore of the separable continent. Entanglement therefore decreases with increasing $h$ but does not experience a sudden death. These features are clearly visible in Fig. 2b. In the following we focus on the case $h=3$, where 2-party and genuine 3-party entanglement are present at zero temperature. For additional data regarding the phase diagram of the model, see SM.
Let us now investigate the fate of entanglement with temperature at $h=3$. From the criterion of the separable ball around the infinite-temperature state [7], we determine the temperatures where $m$-party entanglement is guaranteed to be absent for $m=2,3$. We find $T_{2}=8$ and $T_{3}=22$, respectively. However, these do not tell us if entanglement is absent at lower temperatures. The logarithmic negativity $\mathcal{E}$ and the criterion $W$ vanish at temperatures $T_{\mathcal{E}}=2.1$ and $T_{W}=1.5$, which are much smaller than $T_{2}$ and $T_{3}$. Finally, we observe that for $T_{3}^{*}\approx 3$, the geometric entanglement $\mathcal{D}$ is of order $10^{-4}$, and further rapidly decreases with temperature beyond that point. Hence, the 3-spin density matrix is essentially separable for temperatures much smaller than the rigorous bound $T_{3}$. We report these results in Fig. 2c.
@@
Next, we consider the evolution of entanglement with time during a quantum quench. We initialise the system in a pure product state of up and down spins in the $\sigma^{z}$-basis at $t=0$, and let it evolve under the Ising Hamiltonian discussed above with $h=3$, see SM for details regarding the initial state. In Fig. 2d we show the evolution of our three measures of interest. Both $\mathcal{E}$ and $W$ rapidly rise, reach a maximum, and fall to zero, with $W$ being the shortest lived. We thus see a clear illustration of the sudden death of entanglement for the case of $m=2$ spins. The distance $\mathcal{D}$ follows the same behavior but displays small oscillations close to zero for later times. The state quickly reaches the shallow waters around $t_{3}^{*}\approx 0.4$, while an exact sudden death does not occur for $m=3$, presumably due to the small size of the molecule. An interesting observation is that the non-equilibrium evolution can generate stronger 2- and 3-party entanglement compared to what is found in equilibrium, even at zero temperature, see Fig. 2c. Finally, we observe qualitatively identical behavior for other initial conditions.
@@
As a final example, we consider the evolution of entanglement with separation, working in the groundstate at $h=3$. First, for $m=2$ spins, we find that entanglement disappears when the spins become separated by 2 bonds. Second, for $m=3$ spins, we scale the minimal triangle by $\lambda=2$ so that the sides have bond-length two. In that case we have $W=0$, and we find that the state is extremely close to a separable one, $\mathcal{D}=\mathcal{O}(10^{-6})$, which is many orders of magnitude smaller than for the adjacent triangle.
@@
Figure 2: Icosahedral molecule and its entanglement evolutions. a) Illustration of the quantum molecular magnet with icosahedral geometry. b) Dependence of the geometric entanglement $\mathcal{D}$, the 2-spin logarithmic negativity $\mathcal{E}$ and the genuine 3-party entanglement criterion $W$ in the icosahedral molecule as a function of the transverse field at zero temperature. c) Evolution of the same quantities as a function of temperature with $h=3$. d) Same quantities as a function of time in a quench protocol where the system is prepared in a product state and evolves under the anti-ferromagnetic Ising Hamiltonian with $h=3$.
@@
No continent for fermions.— The above discussion holds for systems of quantum spins. Crucially, spin operators at different sites are independent (they commute with each other) because they represent bosons. However, there exist other particles in nature that do not commute, instead they acquire a minus sign upon exchange: fermions, like electrons or protons. Individual fermion operators possess this relative non-locality with other fermions, but observers nevertheless witness a local world since physical operators are made of an even number of fermions, and are thus bosonic. The density matrix of a physical fermionic system made of $m$ subsystems thus possesses even fermion parity. But how does fermion parity affects the definition of entanglement by constraining viable separable states? A natural notion follows from declaring that a fermionic $m$-party state is un-entangled if it can be written as $\rho^{F}_{\rm sep}=\sum_{k}p_{k}\rho^{(k)}_{1}\otimes\rho^{(k)}_{2}\otimes\cdots\otimes\rho^{(k)}_{m}$, where each $\rho^{(k)}_{j}$ has even parity and the $p_{k}$ form a probability distribution [34, 35, 36]. This definition guarantees that the state has no quantum correlations among the $m$ components, and can thus be prepared locally. Certain states can be brought to this form, but where some $\rho^{(k)}_{j}$do not have even parity. These states cannot be prepared locally and their non-local correlations can be exploited for quantum tasks such as quantum teleportation or quantum data hiding [37, 38, 39].
@@
Given an un-entangled fermionic physical state $\rho^{F}_{\rm sep}$, can we find an entanglement-free region around it as was the case for bosons? We show that the answer is no, and then discuss the consequences. The argument, as given in Ref. [40], is the following. It suffices to consider the case with $m=2$ components by performing a small deformation, $\rho^{F}_{\rm sep}+\delta\rho$, where $\delta\rho$ has zero trace and contains terms of odd fermion parity for subsystem 1. For example, one can hop an odd number of fermions between subsystems 1 and 2, e.g., $\delta\rho=\epsilon(|01\rangle\langle 10|+|10\rangle\langle 01|)$ with positive $\epsilon\ll 1$. The deformed state does not have even parity for subsystem 1, and hence is not separable, even for arbitrarily small $\epsilon$. Therefore, no entanglement-free region exists around a fermionic separable state since nearby states arbitrarily close to $\rho^{F}_{\rm sep}$ contain entanglement between any two subsystems. In the fermionic world, there is thus no separable continent. Separable states lie on a zero-width sand beach, surrounded by an ocean of entangled states arbitrarily close by. We illustrate this in Fig 3.
@@
Turning the tables around, fermionic entanglement is more robust since the state cannot land on a continent upon evolution. Let us say that we heat a fermion system that contains multi-party entanglement at low $T$. The infinite temperature state is un-entangled, but entanglement will decay gradually as a function of temperature without a sudden death. This behavior was observed for integer quantum Hall states [41] and free-fermion systems [34, 42]. Analogous conclusions hold for temporal or spatial evolution. For instance, the entanglement between disjoint regions in fermionic conformal field theories in arbitrary dimensions decays as a power-law with the distance, without a sudden death [29].
@@
Figure 3: Structure of the space of states for fermions. For fermions, because of the fermion-parity superselection rule, there is no separable continent. Instead, separable states form a zero-width sand beach with entangled states arbitrarily close.
Outlook.— In this paper, we discussed how the structure of the space of physical states, and in particular the presence of a separable continent, leads to the sharp disappearance of entanglement in various physical situations, under very generic assumptions. For fermionic systems, the parity superselection rule forbids the existence of the separable continent and drastically modifies the fate of entanglement. In particular, fermionic systems do not experience entanglement sudden death, in stark contrast with bosonic ones. Interestingly, the same conclusion holds for other types of superselection rules, providing a generic recipe for the creation of robust entangled states.
These new insights regarding the structure of entanglement in many-body states will have strong impact for quantum simulation and computation. For instance, efficient algorithms used to simulate quantum matter should incorporate the fact that distant particles are not entangled in spin/boson systems, which strongly constrains the variational space of many-body wave functions. Our work also paves the way for numerous outstanding research avenues. For instance, it will be essential to investigate how multi-party entanglement evolves in a plethora of realistic model Hamiltonians, and to determine the criteria that govern its demise under evolution. We have also seen examples of how quantum matter out-of-equilibrium generates stronger multi-party entanglement than in equilibrium. This is the tip of iceberg: we expect that non-equilibrium dynamics can lead to rich entanglement structures that await to be discovered.
Acknowledgements.— G.P. holds an FRQNT Postdoctoral Fellowship, and acknowledges support from the Mathematical Physics Laboratory of the Centre de Recherches Mathématiques (CRM). W.W.-K. is supported by a grant from the Fondation Courtois, a Chair of the Institut Courtois, a Discovery Grant from NSERC, and a Canada Research Chair.
Supplemental Material: The Fate of Entanglement
Gilles Parez1,2 and William Witczak-Krempa1,2,3
1Département de Physique, Université de Montréal, Montréal, QC H3C 3J7, Canada
2Centre de Recherches Mathématiques,
Université de Montréal, Montréal, QC H3C 3J7, Canada
3Institut Courtois, Université de Montréal, Montréal, QC H2V 0B3, Canada
I A. Entanglement measures and criteria

### v4

@@ IV Icosahedral molecule
The first term corresponds to an anti-ferromagnetic interaction $J>0$ that tends to anti-align neighboring spins along $x$, while the second one is a transverse field that polarises all spins along $z$. Such a molecular quantum magnet is a combination of identical triangular faces, and thus possesses strong geometric frustration. In what follows we shall measure energy in units of the exchange coupling by setting $J=1$. We investigate the fate of 2-, 3- and 4-spin entanglement as a function of magnetic field, temperature, time and separation
IV.1 Full-rank hypothesis and zero temperature phase diagram
As a first step, we checked that the FRH holds for subregions of adjacent spins with 1, 2 and 3 sites by obtaining the $2^{12}$ eigenstates via exact numerical diagonalization, see Appendix B. We then considered the fate of entanglement with varying magnetic field at zero temperature, so that the evolution is parameterized by $s=h$, see Fig. 2b. At $h\!=\!0$, the 2-spin reduced density matrix is a full-rank separable state, which suggests the existence of a sudden-death value $h_{2}^{*}>0$ of the field below which the state is separable, and hence $\mathcal{E}(h<h_{2}^{*})=0$. We find $h_{2}^{*}\approx 0.6$. In contrast, the 3-spin reduced density matrix is not separable at small $h$ and we have $\mathcal{D}>0$ as shown in Fig. 2b. Turning to the GME between the spins on a triangular face, we find that it becomes very small for $h<0.65$, on the order of $10^{-6}$. Using the GTC, we are able to certify that the state is biseparable for $h\leqslant 0.25$. Although the certification becomes more difficult for larger field values, we believe that the true sudden death occurs near $0.65$.
@@ IV.1 Full-rank hypothesis and zero temperature phase diagram
In the opposite limit of $h\to\infty$, the system becomes fully polarized and is in a pure product state. Any $n$-spin density matrix is thus a pure product state, i.e., of rank one, and lies on the boundary of the separable continent. Entanglement for a general $n$-spin density matrix is therefore expected to smoothly decrease with increasing $h$ (at $h\gg 1$) without a sudden death. These features are clearly visible in Fig. 2b for the case of 2- and 3-spin subregions. In the following we focus on the representative case $h=3$, where 2-party and genuine 3-party entanglement are present at zero temperature. For additional data regarding the phase diagram of the model, see Appendix C.
IV.2 Temperature
Let us now investigate the FoE with temperature at $h=3$. From the criterion of the separable ball around the infinite-temperature state [13], we determine the temperatures where $m$-party entanglement is guaranteed to be absent for $m=2,3$. We find $T_{2}=8$ and $T_{3}=22$, respectively. However, these do not tell us if entanglement is absent at lower temperatures. The logarithmic negativity $\mathcal{E}$, as shown in Fig. 2, vanishes at $T_{\mathcal{E}}=1.792$, implying a sudden death of 2-spin entanglement at a much smaller temperature than $T_{2}$. Going beyond 2 spins, with the GTC we can rigorously certify that $T\geqslant T_{3}^{*}=2.5$ leads to full separability of the 3-spin reduced density matrix. This value is close to the point where $\mathcal{D}$ becomes small, see Fig. 2c. GME is expected to disappear before generic multipartite entanglement during the evolution, and indeed we find $T_{W}<T^{*}_{3}$. However, we stress that the condition $W=0$ is not informative, so GME could occur for $T$ up to $T_{3}^{*}$, but certainly not beyond. Going beyond $W$, we have used the generalisation of the GTC to certify the absence of GME beyond a certain temperature, and have obtained an exact upper bound for the death of GME: $T_{\rm GME}^{*}\leqslant 2.0$. In Fig. 2c, the computed values of $\mathcal{D}_{b}$ show a rapid decline of GME, and provide strong evidence that our upper bound is almost tight. To summarize, we have verified that the multipartite 3-spin entanglement, genuine or not, entirely disappears beyond some temperature, in agreement with the general expectation.
IV.3 Quench
@@
Next, we consider the evolution of entanglement with time during a quantum quench. We initialize the system in a pure product state of up and down spins in the $\sigma^{z}$-basis at $t=0$ (in the following, we refer to this quench protocol as Ico-1), and let it evolve under the Ising Hamiltonian discussed above with $h=3$, see Appendix D for details regarding the initial state. In Fig. 2d we show the evolution of various measures, including $\mathcal{D}$ for 4-spin subregion composed of two adjacent triangular faces. All quantities rapidly rise, reach a maximum, and fall to zero, with $\mathcal{D}$ for 4 spins being the longest lived. We thus see a clear illustration of the dynamical sudden death of entanglement for the case of $m=2,3,4$ spins. For 3 spins, we can rigorously and tightly bound the sudden death to occur when $t_{3}^{*}\leqslant 0.369$, while for 4 spins we find $t_{4}^{*}\leqslant 1.5$. To reach a stronger conclusion regarding the fate of GME for 3 spins, we use $\mathcal{D}_{b}$ to obtain a rigorous upper bound on the sudden death time of GME: $t_{\rm GME}^{*}\leqslant 0.25$. Finally, an interesting observation is that the non-equilibrium evolution can generate stronger 2- and 3-party entanglement compared to what is found in equilibrium, even at zero temperature, see Fig. 2c. In order to show that the above behavior is generic, we have studied two other quenches for the icosahedral molecule. Ico-2 corresponds to a time evolution using the $h=3$ Hamiltonian with a distinct initial product state of up and down spins in the $\sigma^{z}$ basis. Ico-3 corresponds to the sudden change of the transverse field, $h=0.04\to 3$; as such, the initial state possesses entanglement (see Fig.2b) in contrast to Ico-1 and Ico-2. For both Ico-2 and Ico-3 protocols, we rigorously show that the 3- and 4-spin states becomes fully separable in finite time, as for Ico-1, see Appendix D.
IV.4 Separation
@@
As a final example, we consider the evolution of entanglement with separation, working in the groundstate at $h=3$. First, for $m=2$ spins, we find that entanglement disappears when the spins become separated by 2 bonds. Second, for $m=3$ spins, we scale the minimal triangle by $\lambda=2$ so that the sides have bond-length two. We find that the state is fully separable using the GTC. We thus see a striking illustration of the short-ranged nature of entanglement, genuine or not, even in a small many-body system.
@@
Figure 2: Icosahedral molecule and its entanglement evolutions. a) Illustration of the quantum molecular magnet with icosahedral geometry. b) Dependence of the geometric entanglement $\mathcal{D}$, the 2-spin logarithmic negativity $\mathcal{E}$ and the genuine 3-party entanglement criterion $\mathcal{D}_{b}$ in the icosahedral molecule as a function of the transverse field at zero temperature. c) Evolution of the same quantities and the criterion $W$ as a function of temperature with $h=3$. d) Same quantities as a function of time in the quench protocol Ico-1, where the system is prepared in a product state and evolves under the anti-ferromagnetic Ising Hamiltonian with $h=3$. In panels b)-c)-d), the open shapes indicate that the corresponding distance measure is rigorously certified to vanish: the state is fully (bi)separable.
V spin chain
We next consider a one-dimensional spin chain with anisotropic nearest-neighbour interactions
$H_{\rm xyz}=\sum_{i=1}^{N}\left(J_{x}\sigma_{i}^{x}\sigma_{i+1}^{x}+J_{y}\sigma_{i}^{y}\sigma_{i+1}^{y}+J_{z}\sigma_{i}^{z}\sigma_{i+1}^{z}-\vec{h}\cdot\vec{\sigma}_{i}\right)$ (7)
with periodic boundary conditions such that $N+1\equiv N$. For our example, we randomly pick in the range $[-1,1]$ the 3 exchange couplings $J_{i}$, and 3 components of the Zeemann magnetic field $h_{i}$: $\vec{J}=(-0.443,0.0938,0.915)$ and $\vec{h}=(0.930,-0.685,0.941)$. We work with $N=14$ sites. The groundstate is unique with energy gap $\Delta E=2.0007$. We have verified that the FRH holds for all $2^{14}$ eigenstates by computing the smallest eigenvalue of the reduced density matrix for a subregion made of 4 adjacent spins, as shown in Appendix B, which confirms the generic nature of $H_{\rm xyz}$.
V.1 Separation
We first examine the FoE with respect to separation in the groundstate. Two spins are separable if they are next nearest neighbours or further apart. For three spins, the GTC certifies the reduced density matrix of sites $(1,3,5)$ to be biseparable, while it certifies full separability for sites $(1,3,6)$. These findings illustrate the short-ranged nature of entanglement in equilibrium.
V.2 Quench
We consider a quench by initializing the chain in a random full product state. Fig. 3 shows the temporal evolution of $\mathcal{D},\mathcal{D}_{b}$ for 3 adjacent sites. Interestingly both quantities oscillate at early times, with $\mathcal{D}$ showing a small revival around $t=5$ before finally becoming zero. Using the GTC, we can certify that for $t\geqslant 5.4$ the subregion becomes fully separable. The GME shows a faster decline, and we can certify biseparability for $t\geqslant 2.4$. We have also been able to show that the subregion of 4 adjacent spins becomes fully separable for $t\geqslant 13$ via the GTC. As above, computing $\mathcal{D}$ is easier than exact certification, and we have obtained $\mathcal{D}(t=9)=3.3\times 10^{-7}$, which strongly suggests that exact separability occurs at earlier times. Finally, to put the non-equilibrium results in context, we give the geometric entanglement of 3 adjacent spins in the groundstate: $\mathcal{D}=0.196$, which is much less than the maximal value achieved during the quench.
Figure 3: Quench for the spin chain. Geometric entanglement $\mathcal{D}$ and $\mathcal{D}_{b}$ as a function of time after a quench in the spin chain defined in Eq. (7). The open shapes indicate that the state is rigorously certified to be fully (bi)separable.
VI A continent without genuine multipartite entanglement for fermions
@@
The above discussion holds for systems of quantum spins. Crucially, spin operators at different sites are independent (they commute with each other) so that they obey bosonic statistics. However, there exist other particles in nature that do not commute, instead they acquire a minus sign upon exchange: fermions, like electrons or protons. Individual fermion operators possess this relative non-locality with other fermions, but observers nevertheless witness a local world since physical operators are made of an even number of fermions, and are thus bosonic. We shall consider fermions hopping on a lattice, as would pertain, for instance, to the fermionic Hubbard model that is used to describe various quantum materials. The density matrix of a physical fermionic system made of $m$ subsystems thus possesses even fermion parity. But how does fermion parity affects the definition of entanglement by constraining viable separable states? A natural notion follows from declaring that a fermionic $m$-party state is un-entangled if it can be written as $\rho^{F}_{\rm sep}=\sum_{k}p_{k}\rho^{(k)}_{1}\otimes\rho^{(k)}_{2}\otimes\cdots\otimes\rho^{(k)}_{m}$, where each $\rho^{(k)}_{j}$ has even parity and the $p_{k}$ form a probability distribution [45, 46, 47]. This definition guarantees that the state has no quantum correlations among the $m$ components, and can thus be prepared locally. Certain states can be brought to this form, but where some $\rho^{(k)}_{j}$do not have even parity. These states cannot be prepared locally and their non-local correlations can be exploited for quantum tasks such as quantum teleportation or quantum data hiding [48, 49, 50]. Moreover, we define biseparable fermionic states as states of the form (2) where each $\rho_{\mathcal{I}_{k}}$ and $\rho_{\overline{\mathcal{I}}_{k}}$ commute with the local fermionic parity of the subsystems they pertain to. Fermionic states thus possess GME if they are not fermionic biseparable.
@@
Given an un-entangled fermionic physical state $\rho^{F}_{\rm sep}$, can we find an entanglement-free region around it as was the case for bosons? We show that the answer is no, and then analyze the same question but for the more interesting case of GME. The argument, adapted from Ref. [51] and generalized to the multipartite case, is the following. Without loss of generality, it suffices to consider the case with $m=2$ components (we can always bipartition the initial multiparty Hilbert space) by performing a small deformation, $\rho^{F}_{\rm sep}+\delta\rho$, where $\delta\rho$ has zero trace and contains terms of odd fermion parity for subsystem 1. For example, one can hop an odd number of fermions between subsystems 1 and 2, e.g., $\delta\rho=\epsilon(|01\rangle\langle 10|+|10\rangle\langle 01|)$ with positive $\epsilon\ll 1$. The deformed state does not have even parity for subsystem 1, and hence is not separable, even for arbitrarily small $\epsilon$. Therefore, no entanglement-free region exists around a fermionic separable state since nearby states arbitrarily close to $\rho^{F}_{\rm sep}$ contain entanglement between any two subsystems. In the fermionic world, there is thus no separable continent. Pictorially, separable states lie on a zero-width sand beach, surrounded by an ocean of entangled states arbitrarily close by. However, we find that entangled states in the vicinity of fermionic separable states are fermionic biseparable by direct construction, see Appendix E. The proof involves taking an un-entangled fermionic physical state, and showing explicitly that a generic small perturbation around it can always be brought into a fermion biseparable form. This implies the existence of a fermionic biseparable continent surrounding the separable beach, and in turn that fermionic systems experience a sudden death of GME during typical evolutions with temperature, time or space. Hence, only non-GME and bipartite entanglement are robust in fermionic systems. We give a schematic illustration of these features in Fig 4.
@@
Turning the tables around, non-GME fermionic entanglement is more robust than in bosonic systems since the state cannot land on a continent upon evolution. Let us say that we heat a fermion system that contains multipartite entanglement at low $T$. The infinite temperature state is un-entangled, but non-GME entanglement will decay gradually as a function of temperature without a sudden death. This behavior was observed for integer quantum Hall states [52] and free-fermion systems [45, 53]. Analogous conclusions hold for temporal or spatial evolution. For instance, the bipartite entanglement between disjoint regions in fermionic quantum critical systems in arbitrary dimensions decays as a power-law with the distance, without a sudden death [39]. In contrast, we predict that any measure of GME in such systems will suffer a sudden death.
@@
Figure 4: Structure of the space of states for fermions. For fermions, because of the fermion-parity superselection rule, there is no separable continent. Instead, separable states form a zero-width sand beach with entangled states arbitrarily close. This beach is surrounded by a light-blue region of biseparable fermionic states, beyond which states possess GME.
VII Outlook
In this paper, we discussed how the structure of the space of physical states, and in particular the presence of a separable continent, leads to the sharp disappearance of entanglement in various physical situations, under very generic assumptions. For fermionic systems, the parity superselection rule forbids the existence of a separable continent and modifies the fate of entanglement. Interestingly, we showed that there is a continent without genuine multipartite fermionic entanglement, which is the richest form of entanglement. So, although fermionic systems do not experience an entanglement sudden death, their genuine multipartite entanglement suffers the same fate as for bosons: it irreversibly disappears during evolution. The same conclusion holds for other types of superselection rules, providing a generic recipe for the creation of robust entangled states.
These new insights regarding the structure of entanglement in many-body states will have strong impact for quantum simulation and computation. For instance, efficient algorithms used to simulate quantum matter should incorporate the fact that distant particles are not entangled in spin/boson systems, which strongly constrains the variational space of many-body wave functions. Our work also paves the way for numerous outstanding research avenues. For instance, it will be essential to investigate how multipartite entanglement evolves in a plethora of realistic model Hamiltonians, and to determine the criteria that govern its demise under evolution. We have also seen examples of how quantum matter out-of-equilibrium generates stronger multipartite entanglement than in equilibrium. This is the tip of iceberg: we expect that non-equilibrium dynamics can lead to rich entanglement structures that await to be discovered.
Acknowledgements
We are grateful to Liuke Lyu for discussions and sharing code. G.P. holds an FRQNT Postdoctoral Fellowship, and acknowledges support from the Mathematical Physics Laboratory of the Centre de Recherches Mathématiques. W.W.-K. is supported by a grant from the Fondation Courtois, a Chair of the Institut Courtois, a Discovery Grant from NSERC, and a Canada Research Chair.
Appendix A Entanglement measures and criteria